OpenStudy (anonymous):

Solve the formula for the indicated letter. Assume that all variables represent nonnegative numbers. A=6s^2, for s (Surface area of a cube with sides of length s)

OpenStudy (radar):

6s^2=A s^2= A/6 \[s=\sqrt{A \over \sqrt{6}}\]

OpenStudy (radar):

\[s=\sqrt{A/6}\]